Monthly loan payments  Personal Finance Problem   Tim Smith is shopping for a used luxury car. He has found one priced at $30,000.

The dealer has told Tim that if he can come up with a down payment of $6,300​, the dealer will finance the balance of the price at a 6​% annual rate over 3 years (36 months).  ​(Hint: Use four decimal places for the monthly interest rate in all your​ calculations.)

a.  Assuming that Tim accepts the​ dealer's offer, what will his monthly​ (end-of-month) payment amount​ be?

b.  Use a financial calculator or spreadsheet to help you figure out what​ Tim's monthly payment would be if the dealer were willing to finance the balance of the car price at an annual rate of 3.8​%?
Answer #1

1.
=(30000-6300)*(6%/12)/(1-1/(1+6%/12)^36)=720.999917601877

2.
=(30000-6300)*(0.3167%)/(1-1/(1+0.3167%)^36)=697.6160493